On this page, we list the main changes that have affected the instrumentand this web site.
- July 28, 2015. The new Giraffe spectral resolutions after the update in the focussing procedure in February have been characterised for all settings. These are the new resolutions:
Setting R_new(Medusa) R_new(IFU/ARGUS)
L385.7 7500 12000
L427.2 6000 8500
L479.7 7500 10150
L543.1 5750 8500
L614.2 6750 10500
L682.2 8500 13500
L773.4 5500 8200
L881.7 6500 10000
H379.0 26750 43350
H395.8 22700 34800
H412.4 31400 48750
H429.7 24000 35550
H447.1A 20250 30250
H447.1B 33000 50530
H465.6 24300 35150
H484.5A 19500 27850
H484.5B 33500 51150
H504.8 23500 37000
H525.8A 18000 28000
H525.8B 31750 50500
H548.8 21500 33700
H572.8 29500 46500
H599.3 20250 32000
H627.3 26400 42200
H651.5A 18000 27900
H651.5B 36900 58350
H665.0 19200 30600
H679.7 21350 35500
H710.5 27400 45500
H737.0A 17350 29250
H737.0B 38100 65100
H769.1 20150 32350
H805.3A 14500 23250
H805.3B 26500 43700
H836.6A 15800 25400
H836.6B 34900 57100
H875.7 18000 28000
H920.5A 11800 17100
H920.5B 23550 34200
- February 27, 2015. FLAMES news appeared on the ESO Science Newsletter (see here). Higher resolution versions of both figures can be found here and here.
- February 9, 2015. On the afternoon of February 2, 2015, an update of the focussing procedure was implemented in FLAMES/Giraffe. This results in significant increase in both spectral resolution and sensitivity across all settings, although these enhencements vary from setting to setting. The Health Check monitor already reflects these variations (for an example, see here). The FLAMES/Giraffe ETC will soon reflect these modifications, once they are well understood at all settings. Users will be contacted by email with further details soon.
- February 9, 2015. The cleaning of the Giraffe HR grating was again successfully performed at the end of January 2015, resulting in an increase in efficiency estimated to be in the range 10-20% for bluer settings.
- December 13, 2014. List of broken fibers updated. After maintenance performed on the fiber positioner during the first week of December, three fibers were recovered. These are Medusa plate 2 RP2, FPS13; IFU sky plate 1, RP153, FPS192; and ARGUS sky, RP5, FPS1. The latter two were enabled before, but had low transmission.
- August 20, 2014. List of broken fibers updated
- May 20, 2014. Update of the ETC with improved efficiency measures at all settings. Estimates of signal-to-noise ratios (SNR) are now more accurate. Compared to the previous ETC version, the updated SNR are generally lower in blue settings and higher in red settings. Significant differences (less than a factor 2) are expected in the most blue and most red settings.
- May 20, 2014. FPOSS manual update. Changes include the addition of two new sections: Known Features and Bugs, and Proper Motion. Also included is a clarification of the field of view for Guide Star selection.
- Dec 18, 2012. Latest attempt to fix fiber 37 yielded 30% of the original transmission. Therefore the fiber will remain disabled until better results are reached or the fiber is replaced. Data obtained with the missing fiber can be reduced with the patched version of the FLAMES/UVES pipeline
- Dec 17, 2012. FLAMES used to test sky-subtraction for the next generation of multi-fibers. Rodrigues et al. (SPIE, 2012) report correction below 1% using cross-beam switching or dual stare method. astro-ph link to be available soon.
- Oct 6, 2012. The UVES fiber 37 on plate 2 is broken since Sep 11. Paranal is working to assess the severity of the problem and the actions to take. In the meanwhile, it was found that the FLAMES-UVES pipeline cannot reduce FLAMES-UVES data taken with a missing fiber. A new release of the pipeline will be made available as soon as possible to allow users to reduce their data. Users willing to re-discuss their observing strategy or to put their OBs on-hold until the problem is solved should contact USD.
- May 11, 2011. GIRAFFE HR cleaned. Lost efficiency in the blue settings is now back.
- May 1, 2011. List of broken fibers updated
- Aug 25, 2010. A new web tool is available to reproduce Figure 16 and Figure 17 of the FLAMES user manual for a specific declination and wavelength.
- Feb 17, 2010. The user manual has been updated with respect to the next P86. A new fast readout mode for GIRAFFE is offered to the community in visitor mode only. The manual can be retrieved at this web page.
- Dec 23, 2009. The current FPOSS version is 1.42. This one is available on the following web page. FPOSS 1.42 must be used for all P85 OBs.
- Dec 7, 2009. The GIRAFFE pipeline was upgraded to 2.8.1 on 2009-09-02. It may be found here.
- Nov 4, 2009. UVES pipeline 4.4.8 supporting new red CCD (and old) data reduction.
- Sep 1, 2009. FLAMES-UVES ETC updated with new QE curve for the upgraded MIT CCD.
- Aug 28, 2009. All manuals updated for P85 including the GIRAFFE reduction cookbook. Please send comments to email@example.com. The main change for P85 is that FLAMES-UVES is more sensitive above 700-nm with less fringing and that the VLT guide star magnitude should now be ideally between mR=9 and 11.
- Aug 26, 2009. The current UVES pipeline will NOT work with the upgraded CCD mosaic. This webpage will be updated when the release has been made.
- Aug 26, 2009. The UVES red arm quantum efficiency ratio for the new/old MIT CCD is shown below (left panel) as derived in the lab and from standard stars. The EEV (bluer part of the red mosaic) is UNCHANGED. The MIT (redder part of the red mosaic) CCD "Nigel" was replaced with another MIT CCD "Zeus" in early July 2009 with superior Quantum Efficiency above 700-nm. The fringing is also much reduced;Plot (a) shows an extracted FLAMES-UVES flatfield order at about 900-nm with the Old CCD counts multiplied by 2.5 (same integration time). Plot(b) shows one normalised order of a stellar spectrum taken before and after the upgrade for standalone UVES.
- Jun 25, 2009. Please remember that the ETC will only be upgraded with the new sensitivity values for the UVES (MIT) red-arm after the recommissioning in early July.When the new values are known they will be posted with a link on this webpage. Until then the current ETC should still be used.
- Jun 23, 2009. Broken and low-throughput fibre page updated for Phase 2. See here for the current status.
- Jun 19, 2009. FLAMES User manual updated for P84 Phase II. Go here for the document.
- Jun 2, 2009. GIRAFFE pipline 2.8 released, see here for details.
- May 1, 2009. Note that because of the upgrade to one of the UVES red CCDs, FLAMES-UVES will not be available until mid-July 2009. FLAMES-GIRAFFE observations continue as normal.
- Feb 25, 2009. Important change for P84 proposals: For P84 for ARGUS only, the GIRAFFE settings L385.7, H379.0, H395.8 and H412.4 will be offered in Visitor Mode ONLY due to the difficulty in obtaining sufficient counts in the attached flatfields. Medusa and IFU are available in Service Mode for all settings. We hope in the future to improve the screen-flat lamp strength in the blue so that we can again offer Argus blue settings in Service Mode.
- Dec 1-3, 2008. FLAMES 6th anniversary meeting; see the ESO Messenger article.The scientific programme with presentations is shown here. An open session at the end of the meeting provided several suggestions for improvements that will be investigated in due course.
- May 25 2008 - Jul 10 2008 and for probably all the old CCD bias images at a lower level: History effect for Biases.The first bias in the normal 5 in the series was sometimes lower by up to about 2 ADU than the subsequent four images. This could cause errors in bias subtraction depending on the parameters used by the GIRAFFE pipeline. All service mode data have been re-reduced using the overscan region and not the master bias to remove this problem. Since2008-07-10 a 2-min gap has been added between the taking of the biases and the effect is no longer present. Note that the effect affected biases taken with the old CCD although at a lower level. See here for details.
- March 1, 2008: The QE of the new detector to the existing one is shown below. A considerable improvement in the sensitivity is expected for wavelengths above 650nm.
- March 16, 2007: Upgrade of GIRAFFE CCD discussed during the last IOT. Upgrade date still to be confirmed (P80 or P81).
- March 03, 2007: FLAMES TCCD successfully replaced by the new generation TCCD.
- April 25, 2005: GIRAFFE CCD decontamination. See details.
- November 23, 2004: Updated user manual for Science Verification and GTO observations
- October 21, 2004: Please find here links to the Science Verification page and the Guarantied Time Observation projects.
- 2004 Dec 7th. Sky flatfields that can be used to create a Solar Spectrum have been taken for all ARGUS HR and LR settings. They will soon be available in the ESO archive. The programme ID used was:60.A-9022(B). The Object name for all data is: SKY_spectrum_field. The dates taken were; 2004-11-26 (LR settings), 2004-11-30 and 2004-12-02(HR settings). January 5, 2005: Updated data for the AO instrument performance.
- 2004 Nov 4th. Sky flatfields that can be used to create a Solar Spectrum have been taken for all Medusa HR and LR settings, plusFLAMES-UVES 520, 580, 860-nm. . They are available in the ESO archive.The programme ID used was: 60.A-9022(B). The Object name for all data is: SKY_spectrum_field. The dates taken were; 2004-09-26, 2004-09-28(HR settings) and 2004-10-11 (LR settings).
- 2004 Sept 7th. Listof GIRAFFE settings which are reduced by the Pipeline on Paranal released.
- 2004 July 29th. Visitor web pages updated.
- 2004 July 19th. New versions of the Manuals released. Main changes are in the User Manual. See here.
- 2004 July 1st. On-going efforts to characterise the velocity accuracy of FLAMES-UVES and FLAMES-GIRAFFE. Please email firstname.lastname@example.org if you want details.
- 2004 June 1. Version of the GIRAFFE pipeline on Paranal has been released since April 2004.
- 2004 April 21. Version of the UVES pipeline released that will reduce both old-and-new format UVES data.See here for details.
- 2004 April 7th P73 preparations. (1) The UVES red format has changed. This means that new FLAMES-UVES data will need to be reduced with the new UVES pipeline. (2) We have added a new column to the "binary FLAMES FIBRE Table" HDU3. This is called "FPD" which stands for "Fibre Position on Detector". It is the last column in the table which now reads "..... 700 870 X Y FPD". FPD is the same as FPS in all cases but ARGUS where the numbering is reversed.(3) We have added a new FITS keyword to the GIRAFFE headers. This is called "INS.EXP.MODE". This matches the "p2pp name" given in the following web pages for the HR grating and the LR grating.
- 2004 March 14th. The approximate efficiencies of the new HR settings have been posted on the Giraffe web page; here.
- 2004 March 4th. From P73 in Visitor Mode (VM) only, and from P74 in both Service Mode (SM) and Visitor mode, seven (7) new HR grating settings will be added. All but one of these work in higher order (therefore higher resolution but lower throughput -sometimes only 25 per cent) than the current settings. A further HR setting now covers both H-alpha and Li6707 (VM only P73, SM/VM P74). See here.
- 2004 January 15th. New versions of the FLAMES user manual and Template reference guide available for P73. Main changes are that (1) there are now 24 HR grating settings as opposed to 22 previously,(2) The OB execution time is calculated automatically within p2pp and (3) ARGUS fast templates are available in p2pp for visitor mode.
- 2003 October: New GIRAFFE higher-efficiency HR grating installed; available for P72. Note that this has lead to a decrease in resolution for some settings. See here for details.
- 2003 July 21: FLAMES User Manual V1.1 released for P72.
- 2003 July: ARGUS mode commissioned successfully; available for P72.For technical details cf. the (public) Commissioning Report
- 2003 June: FACB 149 successfully installed and calibrated on plate 1 and 2
- 2003 May: FLAMES commissioning data released
- 2003 May: Giraffe BLDRS available at SourceForge
- 2003 March: New Pre-Flames Survey Data Release
- 2003 February/March: Early Start of Science Operation Period 71
- 2003 January: Science Verification
- 2002 December: Paranalization II
- 2002 November: Paranalization I
- 2002 October: Commissioning III
- 2002 August: Commissioning II
- 2002 June: Commissioning I
- 2002 April: GIRAFFE installation at UT2
- 2002 March: First Light UVES-fibres
- 2002 February: Installation of OzPoz at UT2
- 2001 December 19: Giraffe First Light in Garching Integration Hall.
- 2001 December: Preliminary Acceptance Australia Fibre Positioner
- 2001 November: Giraffe L-R grating arrived at ESO.
- 2001 October: Commissioning Corrector at Kueyen.
- 2001 September: Acceptance of the Corrector Optics
- 2001 July: FLAMES User Workshop.
- 2001 Jan: status presented at the ESO Wide Review 2001.
- 2001 Jan: Giraffe optical table integrated on the enclosure.
- 2000 Dec: Giraffe collimator and camera completed.
- 2000 July: FLAMES Software FDR successfully passed.
- 2000 May: FLAMES fully under construction.